Find the point of the intersection of the following system of equations algebraically
x^2-Y=5
and 2X+y=3

Answer by checkley71(8403)

X~2-Y=5 THEN Y=-5+X~2 THEN 2X-5+X~2=3 OR X~2+2X-8=0 OR (X+4)(X-2)=0 OR X=-4,2
THEN Y=-5+(-4~2) OR Y=-5+8 OR Y=3 & Y=-5+(2~2) OR Y=-5+4 OR Y=-1